OpenStudy (anonymous):

Which is an equation of a circle with center (5, 0) that passes through the point (1, 1)? a.)(x + 5)2 + y2 = 17 b.)(x - 1)2 + (y - 1)2 = 17 c.)(x - 5)2 + y2 = 17 d.)(x + 5)2 + y2 = 24

OpenStudy (mertsj):

\[(x-h)^2+(y-k)^2=r^2\] That is the equation of a circle with center (h,k) and radius r

OpenStudy (anonymous):

what would be the h and k?

OpenStudy (mertsj):

the center.

OpenStudy (anonymous):

Thankyou!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

(1-5)^2+(1-0)^2= 17^2 17^2= 289 thats not the right answer. did i do something wrong?

OpenStudy (mertsj):

You need to find the radius by using the distance formula to find the distance between the center and (1,1)

OpenStudy (anonymous):

ok